OpenStudy (crashonce):

it is given that x = a^2. If a is decreased by 10% then how much does x decrease by

OpenStudy (crashonce):

@ganeshie8

OpenStudy (mathmath333):

\(x=a^2\) let \(x=100\),then \(a=10\) \(a'=a\times 0.9=10\times 0.9=9\) \(x'=(a')^2=81\) \(\%~decrease=\dfrac{100-81}{100}\times 100=19\) \(\large\tt \begin{align} \color{black}{}\end{align}\)